OpenStudy (4meisu):

Using logs, say N = 4 x 1.356^(0.4t) N = Amount of people. If there are 1200 people in a school, how long would it take for everyone to know a joke? How long would it take for 300 people to know a joke? How many people will know the joke in 1.6 minute?

OpenStudy (4meisu):

I think I know how to solve the first 2 q's but not the one for the 1.6 minute one. Would it be 1200/4 = 1.356^0.4t --> 300 = 1.356^0.4t 0.4t = log1.356 (base)300 0.4t = log300/log1.356 0.4t= 18.72 t= 47 minutes?

OpenStudy (phi):

47 min for how long for 300 people to get the joke looks good. for how many people get the joke in 1.6 minutes, substitute 1.6 for t in the equation and evaluate. you get 4.86 or about 5 people.

OpenStudy (4meisu):

Thank you :)

OpenStudy (4meisu):

But actually, wouldn't the 300 q be 300/4 = 75 so 75 = 1.356^0.4t?